# arXiv:2511.11225v2[cs.IT] 27 Feb 2026

### v2

@@ II-A Mutual Coupling
Mutual coupling arises because the power required to drive a current must work against an electric field generated not only by its own current distribution but also by fields induced from other current distributions. To characterize mutual coupling in CAPAs, let us consider a transmit surface $\mathcal{S}$, which contains a time-harmonic source current to radiate EM waves into free space. In particular, we consider a rectangular planar transmit surface, which is specified by
@@
Let $\bm{j}_{\mathrm{t}}(\mathbf{s})\in\mathbb{C}^{3\times 1},\mathbf{s}\in\mathcal{S}$, denote the Fourier transform of the source current density at frequency $f$, where the corresponding wavelength is $\lambda$. The time factor is assumed to be $e^{-2\pi ft}$. We focus on a narrowband single-carrier system. Wideband multi-carrier systems can be handled by applying the proposed model to each carrier frequency independently.
@@ II-A1 Radiated Power
where the last step is obtained following [11006094, Appendix B].
@@ II-A3 Coupling Kernel
where $\delta(\mathbf{s})$ is the three-dimensional (3D) Dirac delta function.
@@ II-B Uni-Polarized Coupling Kernel
In this paper, we primarily focus on the inter-position coupling and consider a uni-polarized system. Without loss of generality, the polarization direction is assumed to be aligned with the $y$-axis. Under this assumption, the source current is simplified into
@@
Figure 1: Radiation mutual coupling kernel along $x$-axis when $s_{y}=0$.
Figure 2: Radiation mutual coupling kernel along $y$-axis when $s_{x}=0$.
@@ II-C Channel Model
This work focuses on a receiver located in the far-field. Its position is described in spherical coordinates by $\mathbf{r}=R_{0}\mathbf{k}(\theta,\phi)$, where $R_{0}$ is the distance from the origin and $\mathbf{k}(\theta,\phi)=[\cos\theta\sin\phi,\sin\theta\sin\phi,\cos\phi]^{T}$ is the unit direction vector. Under far-field assumptions, the spherical wave represented by the Green’s function can be approximated as a plane wave across the transmit surface as [10220205]
$g(\mathbf{r}-\mathbf{s})=\frac{e^{j\kappa_{0}R_{0}}}{4\pi R_{0}}e^{-j\bm{\kappa}_{\mathrm{r}}^{T}\mathbf{s}},$ (23)
where $\bm{\kappa}_{\mathrm{r}}=\kappa_{0}\mathbf{k}(\theta,\phi)$ is the wavevector in the direction of the receiver. As a consequence, the far-field approximation of the channel response $h(\mathbf{s})$ is given by
$h(\mathbf{s})=\beta e^{-j\bm{\kappa}_{\mathrm{r}}^{T}\mathbf{s}},$ (24)
where $\beta$ is a complex constant that incorporates the path loss and the effect of polarization projection, characterized by
$\beta(\bm{\kappa}_{\mathrm{r}})=\frac{-j\kappa_{0}Z_{0}e^{j\kappa_{0}R_{0}}}{4\pi R_{0}}\left(1-\sin^{2}\theta\sin^{2}\phi\right).$ (25)
@@ III-A Optimal Beamforming
The solution to this problem is characterized by the following theorem.
@@
However, a closed-form expression for $c^{-1}(\mathbf{s})$ is generally unavailable, particularly because the radiation mutual coupling component of the kernel is non-trivial. This makes the direct computation of (30) intractable. In the sequel, a pair of methods are proposed to address this challenge, namely the kernel approximation method and the conjugate method.
@@ III-B1 Wavenumber-domain Approximation
$\mathcal{F}\{f\}(\bm{\kappa})=\iint_{-\infty}^{+\infty}f(\mathbf{s})e^{-j\bm{\kappa}^{T}\mathbf{s}}d\mathbf{s},$ (31)
@@
This expression is derived using the Weyl identity [chew1999waves] and the property that the spatial derivative operator $\partial^{2}_{y}$ corresponds to multiplication by $\kappa_{y}^{2}$ after the Fourier transform. The detailed derivation of (32) is given in Appendix The spatial kernel $c_{\mathrm{rad}}(\mathbf{s})$ is then recovered via the inverse 2D Fourier transform
@@
$=\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}e^{-j\Delta\kappa_{il}^{(x)}s_{x}}ds_{x}\int_{-\frac{L_{y}}{2}}^{\frac{L_{y}}{2}}e^{-j\Delta\kappa_{y,il}s_{y}}ds_{y}$
@@ III-B2 Closed-form Optimal Beamforming
$a_{l}$ $=\int_{\mathcal{S}}h^{*}(\mathbf{z})e^{-j\bm{\kappa}_{l}^{T}\mathbf{z}}d\mathbf{z}=\int_{\mathcal{S}}\beta^{*}e^{-j(\bm{\kappa}_{l}-\bm{\kappa}_{\mathrm{r}})^{T}\mathbf{z}}d\mathbf{z}$
$=\beta^{*}L_{x}L_{y}\mathrm{sinc}\left(\frac{\Delta\widetilde{\kappa}_{x,l}L_{x}}{2}\right)\mathrm{sinc}\left(\frac{\Delta\widetilde{\kappa}_{y,l}L_{y}}{2}\right),$ (44)
with $\Delta\widetilde{\kappa}_{x,l}$ and $\Delta\widetilde{\kappa}_{y,l}$ representing the respective $x$ and $y$ components of $(\bm{\kappa}_{l}-\bm{\kappa}_{\mathrm{r}})$. By substituting (III-B2) into (27), we obtain the closed-form optimal beamformer with mutual coupling as
@@
where $\eta=\int_{\mathcal{S}}|h(\mathbf{s})|^{2}d\mathbf{s}=L_{x}L_{y}|\beta|^{2}$. Finally, the resulting optimal array gain is given by
@@ III-C Conjugate Gradient
Following the principles of the calculus of variations, the optimal solution to (3) is found where the first variation of the functional $\mathcal{J}(v)$, denoted $\delta J(v,\delta v)$, is zero for any perturbation $\delta v$. The first variation is given by
$\delta J(v,\delta v)=\left.\frac{d}{d\epsilon}\mathcal{J}(v+\epsilon\delta v)\right|_{\epsilon=0}=-\Re\left\{\int_{\mathcal{S}}\delta v^{*}(\mathbf{s})r(\mathbf{s})d\mathbf{s}\right\},$ (48)
@@
For the functional to be at a minimum, we must have $\delta J(v,\delta v)=0$ for any arbitrary $\delta v(\mathbf{s})$. This condition implies that the residual $r(\mathbf{s})$ must be zero, which is equivalent to the original Fredholm equation (28). This completes the proof. ∎
@@ III-C2 Numerical Implementation
It is worth noting that for the randomly initialized $\mathbf{v}_{(0)}$, the corresponding residual vector $\mathbf{r}_{(0)}$ need to be calculated directly from its definition in (49), given by
@@
$w_{\mathrm{opt}}(\mathbf{s})=\sqrt{\frac{2P_{\mathrm{t}}}{\mathbf{h}^{T}\mathbf{\Phi}\mathbf{v}^{\mathrm{opt}}}}v_{\mathrm{opt}}(\mathbf{s}).$ (66)
@@
(Applicability to Near-Field Channels) While this work primarily evaluates beamforming under the far-field channel approximation in (24), the proposed optimization frameworks can be discussed in the context of near-field communications. Specifically, the CG method is inherently channel-agnostic and can be directly applied to the exact near-field spherical wave channel model formulated in (II-C). This requires no algorithmic modifications. Conversely, the KA method leverages the far-field plane-wave assumption to analytically evaluate the spatial integrals as a product of $\mathrm{sinc}$ functions, yielding the closed-form solutions in (2) and (III-B2). If applied to the near-field, this specific spatial integral would lose its closed-form tractability, necessitating numerical integration, such as Gauss-Legendre quadrature, to compute the optimal continuous beamformer.
@@ IV-A Array Directivity Analysis
For an infinite transmit aperture; $i.e.,$ $|\mathcal{S}|\rightarrow+\infty$, the optimal array gain with mutual coupling under the far-field assumption is given by
$\lim_{|\mathcal{S}|\rightarrow+\infty}G_{\mathrm{opt}}$ $=\frac{8\pi^{2}|\beta|^{2}}{Z_{s}+C_{\mathrm{rad}}(\widetilde{\bm{\kappa}}_{\mathrm{r}})}=\left(\frac{\kappa_{0}}{R_{0}}\right)^{2}D(\theta,\phi),$ (67)
where $\widetilde{\bm{\kappa}}_{\mathrm{r}}=[\cos\theta\sin\phi,\sin\theta\sin\phi,0]^{T}$ and
@@
The result in Proposition 4 allows us to focus on the impact of polarization and mutual coupling without considering the aperture size. Specifically, if there is no mutual coupling; $i.e.,$ $C_{\mathrm{rad}}(\widetilde{\bm{\kappa}}_{\mathrm{r}})=0$, the directivity of the array gain is merely determined by $|\beta|^{2}$, which is captured by the polarization term $\left(1-\sin^{2}\theta\sin^{2}\phi\right)^{2}$. When mutual coupling exists, the directivity is also determined by the value of the wavenumber-domain mutual coupling kernel at the receiver direction $\widetilde{\bm{\kappa}}_{\mathrm{r}}$. In the following, we focus on the directivity in the E-Plane ($i.e.,$ $y$-$z$ plane aligned with the polarization direction) and the H-Plane ($i.e.,$ $x$-$z$ plane orthogonal to the polarization direction), respectively.
@@
However, a simplified model that omits mutual coupling and assumes the gain is determined solely by the channel response with polarization projection; $i.e.,$ $|\beta|^{2}\propto(1-\sin^{2}\theta\sin^{2}\phi)^{2}$, fails to capture this complete physical picture. Specifically, in the E-plane where $\theta=\pi/2$, this simple model yields a gain proportional to $(1-\sin^{2}\phi)^{2}=\cos^{4}\phi$. This coincidentally predicts that the gain vanishes at the end-fire. In the H-plane where $\theta=0$, the array gain becomes proportional to $(1-0)^{2}=1$. This incorrectly implies the gain is constant for all $\phi$.

### v3

# Mutual Coupling in Continuous Aperture Arrays: Physical Modeling and Beamforming Design

@@ II-A Mutual Coupling
Mutual coupling arises because different current distributions interact through the EM fields they generate, thereby affecting the required driving power and the resulting radiation behavior. To characterize mutual coupling in CAPAs, let us consider a transmit surface $\mathcal{S}$, which contains a time-harmonic source current to radiate EM waves into free space. In particular, we consider a rectangular planar transmit surface, which is specified by
@@
Let $\bm{j}_{\mathrm{t}}(\mathbf{s})\in\mathbb{C}^{3\times 1},\mathbf{s}\in\mathcal{S}$, denote the Fourier transform of the source current density at frequency $f$, where the corresponding wavelength is $\lambda$. The time factor is assumed to be $e^{-j2\pi ft}$. We focus on a narrowband single-carrier system. Wideband multi-carrier systems can be handled by applying the proposed model to each carrier frequency independently.
@@ II-A1 Radiated Power
where the last step is obtained following [11006094, Appendix B]. The real part $\Re\{\cdot\}$ appears because the imaginary part corresponds to reactive, non-radiative field exchange, and therefore does not contribute to the time-averaged radiated power.
@@ II-A3 Coupling Kernel
where $\delta(\mathbf{s})$ is the three-dimensional (3D) Dirac delta function. More particularly, the first term, $Z_{s}\delta(\mathbf{s})\mathbf{I}_{3}$, is a purely local contribution, which models ohmic dissipation at each surface point. In contrast, $\mathbf{G}(\mathbf{s})$ is nonlocal, which captures how currents at one location induce electric fields that affect the power required to sustain currents at other locations.
@@ II-B Uni-Polarized Coupling Kernel
In the sequel, we focus on the inter-position coupling and specialize the general model to a fixed uni-polarized model in order to obtain explicit analytical results. Without loss of generality, the polarization direction is assumed to be aligned with the $y$-axis. The same projection-based method can be applied to any other fixed polarization direction, whereas a genuinely multi-polarized design would require retaining the full matrix-valued coupling model and is beyond the scope of this paper. Under this assumption, the source current is simplified into
@@
Figure 1: Radiation mutual coupling kernel $c_{\mathrm{rad}}(\mathbf{s})$ along the $x$-axis ($s_{y}=0$) at $2.4$ GHz, comparing the proposed polarized model in (15) with the isotropic-kernel model that omits the $\partial_{y}^{2}\varphi(\mathbf{s})$ term. The null locations are only slightly shifted from the conventional half-wavelength rule.
Figure 2: Radiation mutual coupling kernel $c_{\mathrm{rad}}(\mathbf{s})$ along the $y$-axis ($s_{x}=0$) at $2.4$ GHz, comparing the proposed polarized model in (15) with the isotropic-kernel model that omits the $\partial_{y}^{2}\varphi(\mathbf{s})$ term. In the polarization direction, the null locations are significantly shifted and the sidelobes are substantially reduced.
@@ II-C Channel Model
This work focuses on a receiver located in the far-field. For a general far-field direction, define the direction vector $\mathbf{d}(\theta,\phi)=[\cos\theta\sin\phi,\sin\theta\sin\phi,\cos\phi]^{T}$, the associated 3D wavevector $\mathbf{k}=\kappa_{0}\mathbf{d}(\theta,\phi)=[\kappa_{x},\kappa_{y},\kappa_{z}]^{T}$, and its 2D projection onto the transmit-surface plane $\bm{\kappa}=[\kappa_{x},\kappa_{y},0]^{T}$, with $\kappa_{z}(\bm{\kappa})=\sqrt{\kappa_{0}^{2}-\kappa_{x}^{2}-\kappa_{y}^{2}}$. For the fixed receiver direction $(\theta_{0},\phi_{0})$, we further define $\mathbf{k}_{\mathrm{r}}\triangleq\kappa_{0}\mathbf{d}(\theta_{0},\phi_{0})=[\kappa_{\mathrm{r},x},\kappa_{\mathrm{r},y},\kappa_{\mathrm{r},z}]$ and $\bm{\kappa}_{\mathrm{r}}\triangleq[\kappa_{\mathrm{r},x},\kappa_{\mathrm{r},y},0]^{T}$. The receiver position is $\mathbf{r}=R_{0}\mathbf{d}(\theta_{0},\phi_{0})$, where $R_{0}$ is the distance from the origin. Under far-field assumptions, the spherical wave represented by the Green’s function can be approximated as a plane wave across the transmit surface as [10220205]
$g(\mathbf{r}-\mathbf{s})=\frac{e^{j\kappa_{0}R_{0}}}{4\pi R_{0}}e^{-j\mathbf{k}_{\mathrm{r}}^{T}\mathbf{s}}=\frac{e^{j\kappa_{0}R_{0}}}{4\pi R_{0}}e^{-j\bm{\kappa}_{\mathrm{r}}^{T}\mathbf{s}},$ (23)
where we have $\mathbf{k}_{\mathrm{r}}^{T}\mathbf{s}=\bm{\kappa}_{\mathrm{r}}^{T}\mathbf{s}$ since $\mathbf{s}=[s_{x},s_{y},0]^{T}\in\mathcal{S}$. As a consequence, the far-field approximation of the channel response $h(\mathbf{s})$ is given by
$h(\mathbf{s})=\beta_{\mathrm{r}}e^{-j\bm{\kappa}_{\mathrm{r}}^{T}\mathbf{s}},$ (24)
where $\beta_{\mathrm{r}}\triangleq\beta(\theta_{0},\phi_{0},R_{0})$ is a complex constant that incorporates the path loss and the effect of polarization projection, characterized by
$\beta(\theta,\phi,R)=\frac{-j\kappa_{0}Z_{0}e^{j\kappa_{0}R}}{4\pi R}\left(1-\sin^{2}\theta\sin^{2}\phi\right).$ (25)
@@ III-A Optimal Beamforming
The following derivation is written for the scalar beamformer induced by the uni-polarized reduction in Section II. Extending the design to a genuinely multi-polarized system would require a vector-valued functional optimization based on the full matrix-valued coupling kernel $\mathbf{C}(\mathbf{s})$. The solution to this problem is characterized by the following theorem.
@@
However, a closed-form expression for $c^{-1}(\mathbf{s})$ is generally unavailable, particularly because the radiation mutual coupling component of the kernel is non-trivial. This makes the direct computation of (30) intractable. In the sequel, a pair of methods are proposed to address this challenge, namely the kernel approximation method and the conjugate gradient method.
@@ III-B1 Wavenumber-domain Approximation
$\mathcal{F}\{f\}(\bm{\kappa})=\iint_{-\infty}^{+\infty}f(\mathbf{s})e^{-j\bm{\kappa}^{T}\mathbf{s}}ds_{x}ds_{y},$ (31)
@@
This expression is derived using the Weyl identity [chew1999waves] and the property that the spatial derivative operator $\partial^{2}_{y}$ corresponds to multiplication by $\kappa_{y}^{2}$ after the Fourier transform. The detailed derivation of (32) is given in Appendix D. The spatial kernel $c_{\mathrm{rad}}(\mathbf{s})$ is then recovered via the inverse 2D Fourier transform
@@
$=\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}e^{-j\Delta\kappa_{x,il}s_{x}}ds_{x}\int_{-\frac{L_{y}}{2}}^{\frac{L_{y}}{2}}e^{-j\Delta\kappa_{y,il}s_{y}}ds_{y}$
@@ III-B2 Closed-form Optimal Beamforming
$a_{l}$ $=\int_{\mathcal{S}}h^{*}(\mathbf{z})e^{-j\bm{\kappa}_{l}^{T}\mathbf{z}}d\mathbf{z}=\int_{\mathcal{S}}\beta_{\mathrm{r}}^{*}e^{-j(\bm{\kappa}_{l}-\bm{\kappa}_{\mathrm{r}})^{T}\mathbf{z}}d\mathbf{z}$
$=\beta_{\mathrm{r}}^{*}L_{x}L_{y}\mathrm{sinc}\left(\frac{\Delta\widetilde{\kappa}_{x,l}L_{x}}{2}\right)\mathrm{sinc}\left(\frac{\Delta\widetilde{\kappa}_{y,l}L_{y}}{2}\right),$ (44)
with $\Delta\widetilde{\kappa}_{x,l}$ and $\Delta\widetilde{\kappa}_{y,l}$ representing the respective $x$- and $y$-components of $(\bm{\kappa}_{l}-\bm{\kappa}_{\mathrm{r}})$. By substituting (III-B2) into (27), we obtain the closed-form optimal beamformer with mutual coupling as
@@
where $\eta=\int_{\mathcal{S}}|h(\mathbf{s})|^{2}d\mathbf{s}=L_{x}L_{y}|\beta_{\mathrm{r}}|^{2}$. Finally, the resulting optimal array gain is given by
@@ III-C Conjugate Gradient
Following the principles of the calculus of variations, the optimal solution to (3) is found where the first variation of the functional $\mathcal{J}(v)$, denoted $\delta\mathcal{J}(v,\delta v)$, is zero for any perturbation $\delta v$. The first variation is given by
$\delta\mathcal{J}(v,\delta v)=\left.\frac{d}{d\epsilon}\mathcal{J}(v+\epsilon\delta v)\right|_{\epsilon=0}=-\Re\left\{\int_{\mathcal{S}}\delta v^{*}(\mathbf{s})r(\mathbf{s})d\mathbf{s}\right\},$ (48)
@@
For the functional to be at a minimum, we must have $\delta\mathcal{J}(v,\delta v)=0$ for any arbitrary $\delta v(\mathbf{s})$. This condition implies that the residual $r(\mathbf{s})$ must be zero, which is equivalent to the original Fredholm equation (28). This completes the proof. ∎
@@ III-C2 Numerical Implementation
It is worth noting that for an arbitrary initialization $\mathbf{v}_{(0)}$, the corresponding residual vector $\mathbf{r}_{(0)}$ is calculated directly from its definition in (49) as
@@
$w_{\mathrm{opt}}(\mathbf{s})=\sqrt{\frac{2P_{\mathrm{t}}}{\mathbf{h}^{T}\mathbf{\Phi}\mathbf{v}_{\mathrm{opt}}}}v_{\mathrm{opt}}(\mathbf{s}).$ (66)
@@
(Applicability to Near-Field Channels) While this work primarily evaluates beamforming under the far-field channel approximation in (24), the proposed optimization frameworks can be discussed in the context of near-field communications. Specifically, the conjugate gradient method is inherently channel-agnostic and can be directly applied to the exact near-field spherical wave channel model formulated in (II-C). This requires no algorithmic modifications. Conversely, the KA method leverages the far-field plane-wave assumption to analytically evaluate the spatial integrals as a product of $\mathrm{sinc}$ functions, yielding the closed-form solutions in (2) and (III-B2). If applied to the near-field, this specific spatial integral would lose its closed-form tractability, necessitating numerical integration, such as Gauss-Legendre quadrature, to compute the optimal continuous beamformer.
III-C3 Convergence
Finally, we briefly discuss the convergence behavior of the proposed conjugate gradient method. After Gauss-Legendre discretization, solving the continuous optimization problem in (3) can be viewed as solving the following linear system:
$\mathbf{A}\mathbf{v}=\mathbf{h}^{*},\quad\mathbf{A}=\mathbf{C}_{\mathrm{rad}}\mathbf{\Phi}+Z_{s}\mathbf{I},$ (67)
where $\mathbf{\Phi}$ is a diagonal matrix containing the positive quadrature weights. To analyze this system, we introduce the change of variables $\bar{\mathbf{v}}=\mathbf{\Phi}^{1/2}\mathbf{v}$ and $\bar{\mathbf{h}}=\mathbf{\Phi}^{1/2}\mathbf{h}^{*}$, which leads to the following equivalent transformed system:
$\bar{\mathbf{A}}\bar{\mathbf{v}}=\bar{\mathbf{h}},\quad\bar{\mathbf{A}}=\mathbf{\Phi}^{1/2}\mathbf{A}\mathbf{\Phi}^{-1/2}=\mathbf{\Phi}^{1/2}\mathbf{C}_{\mathrm{rad}}\mathbf{\Phi}^{1/2}+Z_{s}\mathbf{I}.$ (68)
Since $\mathbf{\Phi}^{1/2}$ is invertible, the transformed system is fully equivalent to the original one. Moreover, from the wavenumber-domain representation in (32), the radiation kernel has a nonnegative spectrum, which implies that $\mathbf{C}_{\mathrm{rad}}$ is Hermitian positive semidefinite. Hence $\mathbf{\Phi}^{1/2}\mathbf{C}_{\mathrm{rad}}\mathbf{\Phi}^{1/2}$ is also Hermitian positive semidefinite. Since the dissipation term satisfies $Z_{s}>0$, the matrix $\bar{\mathbf{A}}$ is Hermitian positive definite, and the discretized objective is therefore strictly convex with a unique minimizer. It then follows from the classical conjugate gradient theory that the proposed iteration converges to a unique solution for any initialization [shewchuk1994introduction]. In addition, the convergence speed is mainly governed by the condition number of $\bar{\mathbf{A}}$. In practice, a larger quadrature order $M$ increases the problem dimension, and a higher operating frequency makes the coupling kernel more oscillatory. Both effects may worsen the conditioning of the discretized system and therefore slow down the conjugate gradient iterations.
@@ IV-A Array Directivity Analysis
For an infinite transmit aperture, i.e., $|\mathcal{S}|\rightarrow+\infty$, the optimal array gain at the receiver with mutual coupling under the far-field assumption is given by
$\lim_{|\mathcal{S}|\rightarrow+\infty}G_{\mathrm{opt}}$ $=\frac{8\pi^{2}|\beta_{\mathrm{r}}|^{2}}{Z_{s}+C_{\mathrm{rad}}(\bm{\kappa}_{\mathrm{r}})}=\left(\frac{\kappa_{0}}{R_{0}}\right)^{2}D(\theta_{0},\phi_{0}),$ (69)
where
@@
The result in Proposition 4 allows us to focus on the impact of polarization and mutual coupling without considering the aperture size. Specifically, if there is no mutual coupling, i.e., $C_{\mathrm{rad}}(\bm{\kappa}_{\mathrm{r}})=0$, the directivity of the array gain is merely determined by $|\beta_{\mathrm{r}}|^{2}$, which is captured by the polarization term $\left(1-\sin^{2}\theta_{0}\sin^{2}\phi_{0}\right)^{2}$. When mutual coupling exists, the directivity is also determined by the value of the wavenumber-domain mutual coupling kernel at the receiver direction $(\theta_{0},\phi_{0})$. In the following, we focus on the directivity in the E-Plane (i.e., $y$-$z$ plane aligned with the polarization direction) and the H-Plane (i.e., $x$-$z$ plane orthogonal to the polarization direction), respectively.
@@
However, a simplified model that omits mutual coupling and assumes the gain is determined solely by the channel response with polarization projection, i.e., $|\beta|^{2}\propto(1-\sin^{2}\theta\sin^{2}\phi)^{2}$, fails to capture this complete physical picture. Specifically, in the E-plane where $\theta=\pi/2$, this simple model yields a gain proportional to $(1-\sin^{2}\phi)^{2}=\cos^{4}\phi$. This coincidentally predicts that the gain vanishes at the end-fire. In the H-plane where $\theta=0$, the array gain becomes proportional to $(1-0)^{2}=1$. This incorrectly implies the gain is constant for all $\phi$.